# The Impacts of the Gender Imbalance on the Marriage Market:

### v3

@@ C.3 Endogeneity Test under Control Function Approach
  • (2) Akita War-bereaved Family Association. Yonjyūnen shi: Akita-ken Izoku Rengōkai yonjyūnen kinen shi, (Akita War-bereaved Family Association 40th Anniversary). [in Japanese]. Akita War-bereaved Family Association, Akita, 1994.
@@
  • (4) Bureau of Statistics, Office of the Prime Minister. Syōwanijyuḡonen Kokuseichōsahōkoku, dainanakan, fukenhen (1950 Population Census of Japan, prefecture part (volume 7)). [in Japanese]. Bureau of Statistics, Office of the Prime Minister, Tokyo, 1951a.
@@
  • (8) Bureau of Statistics, Office of the Prime Minister. Syōwasanjyun̄en Kokuseichōsahōkoku, daigokan, fukenhen (1955 Population Census of Japan, prefecture part (volume 5)). [in Japanese]. Bureau of Statistics, Office of the Prime Minister, Tokyo, 1956a.
@@
  • (14) Fukui War-bereaved Family Association. Fukui-ken izoku-rengōkai yonjyūnen no ayumi, (Fukui War-bereaved Family Association 40th Anniversary History). [in Japanese]. Fukui War-bereaved Family Association, Fukui, 1988.
@@
  • (16) Fukushima War-bereaved Family Association. Fukushima-ken izoku-kai no ayumi: Fukushima-ken izoku-kai gojyūnen shi, (Fukushima War-bereaved Family Association 50th Anniversary History). [in Japanese]. Fukushima War-bereaved Family Association, Fukushima, 1999.
@@
  • (18) Gunma Prefecture. Gunma-ken fukuin engo shi, (Gunma Demobilization Support History). [in Japanese]. Gunma Prefecture, Gunma, 1974.
@@
  • (20) Harald Fuess. Meijiki no konin to rikon ibunkatekishiten de mita kindai syakai no seni, (Marriage and Divorce in the Meiji Era: Transitions in Modern Society from a Cross- Cultural Perspective). [in Japanese]. In Satomi Kurosu, editor, Rekishijinkōgaku kara mita kekkon, rikon, saikon, (Marriage, Divorce, and Remarriage from the Perspective of the Historial Demography). [in Japanese], pages 171–203. Reitaku University Press, Chiba, 2012.
@@
  • (22) Investigation Department for Unrepatriated Persons. Sōsyūtsu Todōfuken betsu Tandoku Kaitakudan tō Ichiranhyō, (the List of Manchuria Pioneer Groups by Prefecture). [in Japanese]. Tokyo, 1954
@@
  • (26) Tomoko Kanzaki. Sengo nihon jyosei seiakushi, (A History of post-war Japanese Women’s Policy). [in Japanese]. Akashi Shoten, Tokyo, 2009.
@@
  • (28) Noriko Kurushima, Hiroko Nagano, and Shizue Osa. Jendā kara mita nihonshi, (Japanese History from a Gender Perspective). [in Japanese]. Otsuki Shoten, Tokyo, 2015.
@@
  • (32) Nagasaki Prefecture. Engo no ayumi: Syūsen 30 shūnen kinen, (Nagasaki Support History: 30th Anniversary after the War). [in Japanese]. Nagasaki Prefecture, Nagasaki, 1976.
@@
  • (34) National Institute of Population and Social Security Research. Shusshō Dōkō Chōsa: Nihonjin no Kekkon to Syussan (Basic Survey on Fertility Trends: Marriage and Child-birth in Japan). [in Japanese]. National Institute of Population and Social Security Research, Tokyo, 1998.
@@
  • (36) Masakatsu Okado. 1950 Nendai no Nōminkazoku : Jisannan to Jyosei no Ishiki wo Chus̄hin ni, (Peasant families in the 1950s: Focusing on the attitudes of the second and third sons and women). [in Japanese]. In Masanori Nakamura, editor, Kingendai Nihon no Shinshiten: Keizaishi kara no Apurc̄hi, (New Perspectives on Modern and Con- temporary Japan: An Approach from Economic History). [in Japanese], pages 57–81. Yoshikawakōbunkan, Tokyo, 2000.
@@
  • (38) Okayama Prefecture. Engo no ayumi: Sengo gojyū shūnen kinen, (Okayama Support History: 50th Anniversary after the War). [in Japanese]. Okayama Prefecture, Okayama, 1996.
@@
  • (40) Shiga War-bereaved Family Association. Shiga-ken izoku-kai no ayumi, (Shiga War-bereaved Family Association History). [in Japanese]. Shiga War-bereaved Family Association, Shiga, 1995.
@@
  • (42) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, fukenhen (1930 Popu- lation Census of Japan, prefecture part). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1933.
@@
  • (44) Statistics Bureau of the Cabinet. Shōwajun̄en Kokuseichōsahōkoku, fukenhen (1935 Popu- lation Census of Japan, prefecture part). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1939a.
@@
  • (46) Takao Terawaki. Sengo shori, Ikazoku engo, Fujin hogo : Shiryōsyū sengo nihon no syakai fukushi seido, (Collection of Materials of Postwar Social Welfare System in Japan). [in Japanese], Kashiwa shobō, Tokyo, 2015.
@@
  • (48) Toyama Prefecture. Toyama-ken shūsen shori shi, (Toyama History after the War). [in Japanese]. Toyama Prefecture, Toyama, 1975.
@@
  • (52) Jeffrey M. Wooldridge. 2015. Control Function Methods in Applied Econometrics. The Journal of Human Resources vol. 50, no. 2, pp. 420–445.
@@
  • (54) Haruko Wakita, Reiko Hayashi, and Kazuko Nagahara. Nihon jyosei shi, (Japanese Women’s History). [in Japanese]. Yoshikawa Kobunkan, Tokyo, 2011.
@@
  • (56) Yamagata War-bereaved Family Association. Yamagata-ken izoku-kai gojyūnen shi, (Yamagata War-bereaved Family Association 50th Anniversary). [in Japanese]. Yamagata War-bereaved Family Association, Yamagata, 1995.
@@
  • (58) Yoshihiko Yuzawa. Sengo kazoku hendō no tōkeiteki kōsatu, (A Statistical Study of Post- war Family Change). [in Japanese]. The University of Tokyo Press, Tokyo, 1977.

### v4

@@ C.3 Endogeneity Test under Control Function Approach
C.4 Reduced Form Results
Table C.2 summarizes the results from the reduced-form regression. Panels A-1 and A-2 show the results for women and men in the 1950 sample, whereas Panels B-1 and B-2 show the results for women and men in the 1955 sample. Unline the estimates obtained from out instrumental variable approach, the estimates shown in this table are essentially biased due to the endogeneity. As shown, the estimates from the reduced-form regression are not stable. For example, while the results for women show similar estimates both in the 1950 and 1955 samples, the magnitudes are systematically smaller than those from the instrumental variable estimation (Panels A-1 and B-1 in Table 2). Moreover, the results for men differ between the 1950 and 1955 samples, while the results from the instrumental variable estimation remain quite stable. (Panels A-2 and B-2 in Table 2).
Table C.2: Reduced Form Results:
1950 and 1955 Population Census Data
Dependent Variables
(1) Single (2) Married (3) Divorced (4) Widowhood
Panel A: 1950 Population Census
      Panel A-1: Women
      Sex Ratio $-91.22$ $249.63$ 30.46-30.46***- 30.46 ∗ ∗ ∗ 160.63-160.63***- 160.63 ∗ ∗ ∗
$(111.00)$ $(185.58)$ $(10.04)$ $(28.50)$
      Age fixed-effect Yes Yes Yes Yes
      Control variables Yes Yes Yes Yes
      Mean of the dependent variables $194.26$ $676.88$ $22.44$ $87.23$
      SD of the dependent variables $297.81$ $260.53$ $9.85$ $69.49$
      Observations 1,564 1,564 1,564 1,564
      Panel A-2: Men
      Sex Ratio $46.04$ $-70.29$ 13.85-13.85***- 13.85 ∗ ∗ ∗ $-4.82$
$(126.28)$ $(177.13)$ $(4.34)$ $(4.81)$
      Age fixed-effect Yes Yes Yes Yes
      Control variables Yes Yes Yes Yes
      Mean of the dependent variables $263.70$ $693.98$ $9.94$ $13.00$
      SD of the dependent variables $357.02$ $355.63$ $5.23$ $15.52$
      Observations 1,564 1,564 1,564 1,564
Panel B: 1955 Population Census
      Panel B-1: Women
      Sex Ratio $71.59$ $104.17$ 37.06-37.06***- 37.06 ∗ ∗ ∗ 140.01-140.01***- 140.01 ∗ ∗ ∗
$(76.93)$ $(92.98)$ $(11.13)$ $(19.32)$
      Age fixed-effect Yes Yes Yes Yes
      Control variables Yes Yes Yes Yes
      Mean of the dependent variables $228.47$ $672.19$ $23.79$ $75.53$
      SD of the dependent variables $321.78$ $269.88$ $12.32$ $73.95$
      Observations 1,564 1,564 1,564 1,564
      Panel B-2: Men
      Sex Ratio 121.48121.48***121.48 ∗ ∗ ∗ 100.04-100.04**- 100.04 ∗ ∗ 16.03-16.03***- 16.03 ∗ ∗ ∗ $-4.84$
$(43.89)$ $(46.12)$ $(5.73)$ $(2.89)$
      Age fixed-effect Yes Yes Yes Yes
      Control variables Yes Yes Yes Yes
      Mean of the dependent variables $290.83$ $690.38$ $10.25$ $8.43$
      SD of the dependent variables $379.99$ $369.32$ $6.07$ $10.75$
      Observations 1,564 1,564 1,564 1,564
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors from the cluster-robust variance estimation reported in parentheses are clustered at the 46-prefecture level.
Notes: This table shows the OLS estimates for equation 2. The dependent variables used in columns (1)–(4) of Panel A (Panel B) are the proportion of single, married, divorced, and widowed women (men) per 1,000 women (men), respectively. All the regressions include the share of workers employed in munitions factories, the mortality rate of homefront people due to air bombings, the Manchurian settlers’ mortality, and age-fixed effects. The number of observations is $1,564$ (46 prefectures $\times$ 34-age range) in all the regressions.
@@ C.4 Reduced Form Results
  • (2) Akita War-Bereaved Family Association. Yonjyūnen shi: Akita-ken Izoku Rengōkai yonjyūnen kinen shi (Akita Warbereaved Family Association, 40th anniversary). [In Japanese]. Akita War-bereaved Family Association, Akita, 1994.
@@
  • (4) Bureau of Statistics, Office of the Prime Minister. Syōwanijyuḡonen Kokuseichōsahōkoku, dainanakan, fukenhen (1950 Population Census of Japan, prefecture (Volume 7)). [In Japanese]. Bureau of Statistics, Office of the Prime Minister, Tokyo, 1951a.
@@
  • (8) Bureau of Statistics, Office of the Prime Minister. Syōwasanjyun̄en Kokuseichōsahōkoku, daigokan, fukenhen (1955 Population Census of Japan, prefecture (Volume 5)). [In Japanese]. Bureau of Statistics, Office of the Prime Minister, Tokyo, 1956a.
@@
  • (14) Fukui War-bereaved Family Association. Fukui-ken izoku-rengokai yonjyunen no ayumi, (Fukui War-bereaved Family Association 40th Anniversary History). [In Japanese]. Fukui Warbereaved Family Association Fukui, 1988.
@@
  • (16) Fukushima War-bereaved Family Association. Fukushima-ken izoku-kai no ayumi: Fukushima-ken izoku-kai gojyūnen shi (Fukushima War-Bereaved Family Association 50th anniversary). [In Japanese]. Fukushima War-Bereaved Family Association, Fukushima, 1999.
@@
  • (18) Gunma Prefecture. Gunma-ken fukuin engo shi (Gunma demobilization support history). [In Japanese]. Gunma Prefecture Gunma. 1974.
@@
  • (20) Harald Fuess. Meijiki no konin to rikon ibunkatekishiten de mita kindai syakai no seni (Marriage and divoce in the Meiji Era: Transitions in modern society from a cross-cultural perspective). [In Japanese]. In Satomi Kurosu, editor Rekishijinkōgaku kara mita kekkon, rikon, saikon (marriage, divorce, and re-arriage from the perspective of Historial demography). [In Japanese], pages 171–203. Reitaku University Press, Chiba (2012).
@@
  • (22) Investigation Department for Unrepatriated Individuals. Sōsyūtsu Todōfuken betsu Tandoku Kaitakudan tō Ichiranhyō ( List of Manchurian Pioneer Groups by Prefecture). [In Japanese]. Tokyo, 1954
@@
  • (26) Tomoko Kanzaki. Sengo nihon jyosei seiakushi (A History of post-war Japanese Women’s policy). [In Japanese]. Akashi Shoten, Tokyo, 2009.
@@
  • (28) Noriko Kurushima, Hiroko Nagano, and Shizue Osa. Jendk̄ara mita nihonshi (Japanese History from a Gender Perspective). [In Japanese]. Otsuki Shoten, Tokyo, 2015.
@@
  • (32) Nagasaki Prefecture. Engo no ayumi: Syūsen 30 shūnen kinen (Nagasaki Support History: 30th Anniversary after the War). [In Japanese]. Nagasaki Prefecture, Nagasaki, 1976.
@@
  • (34) National Institute of Population and Social Security Research. Shusshō Dōkō Chōsa: Nihonjin no Kekkon to Syussan (basic survey on fertility trends: Marriage and childbirth in Japan). [In Japanese]. National Institute of Population and Social Security Research, Tokyo, 1998.
@@
  • (36) Masakatsu Okado. 1950 Nendai no Nōminkazoku: Jisannan to Jyosei no Ishiki wo Chus̄hin ni, (Peasant families in the 1950s: Focusing on the attitudes of the second and third sons and women). [In Japanese]. In Masanori Nakamura, editor, Kingendai Nihon no Shinshiten: Keizaishi kara no Apurc̄hi (New perspectives on modern and contemporary Japan: An Approach from Economic History). [In Japanese], pages 57–81. Yoshikawak: obunkan, Tokyo, 2000.
@@
  • (38) Okayama Prefecture. Engo no ayumi: Sengo gojyū shūnen kinen (Okayama Support History: 50th Anniversary after the War). [In Japanese]. Okayama Prefecture, Okayama, 1996.
@@
  • (40) Shiga War-bereaved Family Association. Shiga-ken izoku-kai no ayumi, (Shiga warbereaved family association history). [In Japanese]. Shiga War-Bereaved Family Association Shiga. 1995.
@@
  • (42) Statistics Bureau of the Cabinet. Shōwagonen, Kokuseichōsahōkoku, fukenhen (1930 Popu- lation Census of Japan, prefecture section). [In Japanese]. Statistics Bureau of the Cabinet. Tokyo, 1933.
@@
  • (44) Statistics Bureau of the Cabinet. Shōwajun̄en, Kokuseichōsahōkoku, fukenhen (1935 Popu- lation Census of Japan, Prefecture). [In Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1939a.
@@
  • (46) Takao Terawaki. Sengo shori, Ikazoku engo, Fujin hogo: Shiryōsyū sengo nihon no syakai fukushi seido (A Collection of Materials of Postwar Social Welfare System in Japan). [In Japanese] Kashiwa shobō, Tokyo, 2015.
@@
  • (48) Toyama Prefecture. Toyama-ken shūsen shori shi (Toyama history after the war). [In Japanese]. Toyama Prefecture, Toyama 1975.
@@
  • (52) Jeffrey M. Wooldridge. 2015. Control-function methods in applied econometrics. Journal of Human Resources vol. 50, no. 2, pp. 420–445.
@@
  • (54) Haruko Wakita, Reiko Hayashi, Kazuko Nagahara. Nihon jyosei shi (Japanese Women’s history). [In Japanese]. Yoshikawa Kobunkan, Tokyo, 2011.
@@
  • (56) Yamagata War-Bereaved Family Association. Yamagata-ken izoku-kai gojyūnen shi (Yamagata Warbereaved Family Association, 50th anniversary). [In Japanese]. Yamagata War-bereaved Family Association, Yamagata, 1995.
@@
  • (58) Yoshihiko Yuzawa. Sengo kazoku hendō no tōkeiteki kōsatu (A Statistical Study of Postwar Family Change). [In Japanese]. Tokyo: University of Tokyo Press, Tokyo, 1977.